OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? Find the exact total surface area of a cone that has a slant height of 24 inches and a radius of 3 inches. a. 678.24 in.2 b. 254.34 in.2 c. 508.68 in.2 d. 226.08 in.2

OpenStudy (anonymous):

@sleepyjess please help

OpenStudy (sleepyjess):

 that is slant height correct?

OpenStudy (anonymous):

@sleepyjess yes...

OpenStudy (sleepyjess):

ok just making sure \(A=\pi r(r+\sqrt{h^2+r^2)}\)

OpenStudy (sleepyjess):

but to use that formula we need the height, not the slant

OpenStudy (sleepyjess):

to find h we use \(\sqrt{l^2-r^2}\)

OpenStudy (anonymous):

so 24.2 times 3.2

OpenStudy (sleepyjess):

no (24*24)-(3*3)

OpenStudy (anonymous):

@sleepyjess thats what i meant .

OpenStudy (sleepyjess):

ok then yes lol

OpenStudy (anonymous):

@sleepyjess i got 567

OpenStudy (sleepyjess):

ok and the square root of 567 is 23.8

OpenStudy (anonymous):

@sleepyjess can you post your report on my question please?

OpenStudy (anonymous):

@sleepyjess what would i do from there ?

OpenStudy (sleepyjess):

yeah give me 5 min? @Monster1234

OpenStudy (anonymous):

Omg thank you!

OpenStudy (sleepyjess):

now use \(A=\pi r(r+\sqrt{h^2+r^2)}\) using 3 for r and 23.8 for h

OpenStudy (anonymous):

@sleepyjess i got 2017.7 ?

OpenStudy (sleepyjess):

hmmm i got approx 311 i rounded most of my stuff but it still should have come close to one of the answers lets see what went wrong \(\sqrt{23.8^2+3^2}~~~~~\sqrt{566.44+9}~~~~~~\sqrt{575.44}~~~~~~23.9883304963\) now we have \(\pi 3(3+23.9883304963)~~~~~~\pi 3(26.9883304963)\) \(\pi 3=9.42\) \(9.42(26.9883304963)=253.96018997\)

OpenStudy (sleepyjess):

with some things slightly rounded so B is probably the answer they are looking for here